THE RATIO OF MALES TO FEMALES AT John High School is 5:4. If there are 800 females that attend Jones, how many males attend?
Answer by stanbon(75887)

THE RATIO OF MALES TO FEMALES AT John High School is 5:4. If there are 800 females that attend Jones, how many males attend?
----
5:4 is the same as 5x:4x where 5x is # of males and 4x is # of females.
----
Equation:
5x + 4x = 800
9x = 800
x = 800/9
---
# of males = 5(800/9) = 4000/9 = 444 when rounded down to a whole number
======================
Cheers,
Stan H.